# Mixed precision iterative refinement for least squares with linear equality constraints and generalized least squares problems

### v2

@@ 5.3 Block-diagonal split preconditioner for the GLS problem
$\begin{split}M_{l}&=\operatorname{diag}\bigl{\{}\alpha^{-1/2}I_{p},\alpha^{1/2}\hat{T}_{2}^{-1}Q^{*},\alpha^{-1/2}S^{*}\hat{R}^{-*}\bigr{\}},\\ M_{r}&=\operatorname{diag}\bigl{\{}\alpha^{-1/2}I_{p},\alpha^{1/2}Q\hat{T}_{2}^{-*},\alpha^{-1/2}\hat{R}^{-1}S\bigr{\}},\end{split}$ (5.5)
@@
where $\check{I}_{m}=[I_{m},0]^{*}\in\mathbb{C}^{n\times m}$; $Z=[Z_{1}^{*},Z_{2}^{*}]^{*}$ with $Z_{2}\in\mathbb{C}^{p\times n}$. Note that the approximation pertains solely to the rounding error similar to the arguments in Section 5.2; in exact arithmetic (5.6) is satisfied as an equality.
@@
$\begin{split}M_{l}&=\operatorname{diag}\bigl{\{}\alpha^{-1/2}I_{p},\alpha^{1/2}U^{-1}Q^{*},\alpha^{-1/2}Y^{*}\hat{R}^{-*}\bigr{\}},\\ M_{r}&=\operatorname{diag}\bigl{\{}\alpha^{-1/2}I_{p},\alpha^{1/2}QU^{-*},\alpha^{-1/2}\hat{R}^{-1}Y\bigr{\}},\end{split}$ (5.7)
@@
$\kappa_{\infty}(M_{l}\tilde{F}M_{r})\leq C\cdot(m+n+p).$
The main objective of the GMRES solver is to overcome the limitation on the condition number $\kappa_{\infty}(\tilde{F})$ in Section 4. Compared to traditional iterative refinement techniques that have a convergence condition number limit of $\mathcal{O}(\bm{u}_{f}^{-1})$, it has been showed in [10] that GMRES algorithms, if implemented carefully, can relax this constraint to $\mathcal{O}(\bm{u}_{f}^{-1}\bm{u}^{-1/2})$. The limiting accuracy of the mixed precision GMRES algorithm is no different to the classical iterative refinement ones in Theorems 4.8 and 4.10. A specific example of that can be found in Table 3 by employing IEEE half/single/double precisions.
@@
Method $\bm{u}_{f}$ $\bm{u}$ $\bm{u}_{r}$ Upper bound on $\kappa_{\infty}(\tilde{F})$ Forward error
@@ 6 Numerical experiments
In this section, we present numerical results of our mixed precision LSE and GLS algorithms. In our tests, we set the combinations of precisions as $(\bm{u}_{f},\bm{u}_{s},\bm{u},\bm{u}_{r})=(\text{single},\text{single},\text{double},\text{double})$ in IEEE standard. All experiments are performed on a Linux server equipped with two sixteen-core Intel Xeon Gold 6226R 2.90 GHz CPUs with 1024 GB of main memory.
@@ 6.1 Experiment settings
    DGGLSE [the reference algorithm]: the LAPACK subroutine using the null-space approach via the GRQ factorization performed totally in double precision.
@@
    MPLSE-GMRES-Left: mixed precision LSE algorithm using GMRES algorithm with the left preconditioner (5.1) as iterative refinement.
@@
In our implementation, the GMRES process is conducted in double precision entirely. Carson and Daužickaitė discussed using lower precision in GMRES-based iterative refinement in [8]. However, for dense matrices, there is no need for that, as the most time-consuming part, matrix–vector multiplication, needs to be performed in the working precision regardless. We choose $\alpha=\lVert r_{0}\rVert_{2}$ and $\alpha=\lVert y_{0}\rVert_{2}$ in the GMRES preconditioner for LSE and GLS problems, respectively. The parameter tol is set to $10^{-13}$ for the stopping criteria of iterative refinement.
@@ 6.2 Tests for the LSE problem
For the LSE problem, cases $m>n\gg p$ and $m\gg n>p$ are often of interest in practice. In case $m>n\gg p$, we choose $n\in\{1024,2048,3072\}$ in our experiments, and set $(m,p)=(8n,n/32)$, $(8n,n/64)$, $(10n,n/32)$, $(10n,n/64)$, $(12n,n/32)$, and $(12n,n/64)$ to generate different test matrices with fixed condition number $\kappa_{2}\bigl{(}[A^{*},B^{*}]\bigr{)}=10^{3}$, $10^{5}$, $10^{7}$, respectively. In case $m\gg n>p$, we choose $n\in\{128,192,256\}$, and set $(m,p)=(128n,n/8)$, $(128n,n/4)$, $(128n,n/2)$, $(256n,n/8)$, $(256n,n/4)$, and $(256n,n/2)$ with fixed condition number $\kappa_{2}\bigl{(}[A^{*},B^{*}]\bigr{)}=10^{5}$. The matrix $[A^{*},B^{*}]$ with a specified condition number is generated through $[A^{*},B^{*}]=W_{1}\Sigma W_{2}$, where $\Sigma$ is a diagonal matrix with the given condition number produced by DLATM1 from LAPACK, and $W_{1}$ and $W_{2}$ are the unitary factors from the QR factorization of randomly generated matrices.
@@ 6.2.1 Case $m>n\gg p$
MPLSE err-1 $3.3\cdot 10^{-17}$ $2.0\cdot 10^{-16}$ $2.2\cdot 10^{-14}$ $2.6\cdot 10^{-10}$
@@
Figure 1: Relative run time of MPLSE, MPLSE-GMRES-Left, and MPLSE-GMRES-BD compared to DGGLSE for matrices with $\kappa_{2}\big{(}[A^{*},B^{*}]\big{)}=10^{3}$. For each matrix, the three columns from left to right represent the results of MPLSE, MPLSE-GMRES-Left, and MPLSE-GMRES-BD, respectively.
Figure 2: Relative run time of MPLSE, MPLSE-GMRES-Left, and MPLSE-GMRES-BD compared to DGGLSE for matrices with $\kappa_{2}\big{(}[A^{*},B^{*}]\big{)}=10^{5}$. For each matrix, the three columns from left to right represent the results of MPLSE, MPLSE-GMRES-Left, and MPLSE-GMRES-BD, respectively.
Figure 3: Relative run time of MPLSE, MPLSE-GMRES-Left, and MPLSE-GMRES-BD compared to DGGLSE for matrices with $\kappa_{2}\big{(}[A^{*},B^{*}]\big{)}=10^{7}$. For each matrix, the three columns from left to right represent the results of MPLSE, MPLSE-GMRES-Left, and MPLSE-GMRES-BD, respectively.
From Table 4, Figures 13, we observe that in comparison with the left preconditioner, the GMRES-based algorithm with the block-diagonal split preconditioner requires more iterations to converge. However, the left preconditioner costs more execution time due to the higher computational cost of applying (5.1) to a vector. This indicates that it is preferable to use the block-diagonal split preconditioner (5.3) or (5.4) for GMRES-based iterative refinement.
@@
As the condition number $\kappa_{2}\bigl{(}[A^{*},B^{*}]\bigr{)}$ becomes $10^{7}$, which is close to $\bm{u}_{f}^{-1}$, all three mixed precision LSE algorithms require more iterations and time to converge; see Figure 3. This is due to the reduced accuracy of the initial guess and GRQ factorization, which also affects the quality of the preconditioner. Our analysis indicates that MPLSE is not suitable for scenarios where condition numbers exceed approximately $\bm{u}_{f}^{-1}$. In Figure 3, we see that MPLSE-GMRES-BD remains approximately $10\%$$20\%$ faster than DGGLSE, a speedup between 1.11 and 1.25. However, that will diminish rapidly as the condition number increases further.
@@ 6.2.2 Case $m\gg n>p$
Figure 4: Relative run time of MPLSE, MPLSE-GMRES-Left, and MPLSE-GMRES-BD compared to DGGLSE for matrices with $\kappa_{2}\big{(}[A^{*},B^{*}]\big{)}=10^{5}$. For each matrix, the three columns from left to right represent the results of MPLSE, MPLSE-GMRES-Left, and MPLSE-GMRES-BD, respectively.
@@ 6.3 Tests for the GLS problem
For the GLS problem, we choose the dimension $n\in\{1024,2048,3072\}$ in our experiments, and set $(p,m)=(8n,n/32)$, $(8n,n/64)$, $(10n,n/32)$, $(10n,n/64)$, $(12n,n/32)$, and $(12n,n/64)$ to generate different test matrices with fixed condition number $\kappa_{2}\big{(}[W,V]\big{)}=10^{3}$, $10^{5}$, $10^{7}$, respectively. Note that we do not test GMRES with the left preconditioner (5.2) for the GLS problem, since it has already been shown in Section 6.2 that it is preferable to use the block-diagonal split preconditioner.
@@
Table 5 presents the accuracy of the results and the iteration counts for a single test matrix of consistent size across varying condition numbers. In Figures 57, we present the relative run time for matrices with fixed condition numbers $\kappa_{2}\bigl{(}[W,V]\bigr{)}=10^{3}$, $10^{5}$, $10^{7}$. For test matrices with relatively small condition numbers, both algorithms reduce the computational time by a factor around $40\%$$50\%$, i.e., a speedup between $1.66$ and $2.00$, with MPGLS being slightly faster. The analysis shows that MPGLS fails when condition numbers exceed approximately $\bm{u}_{f}^{-1}$. MPLSE-GMRES-BD can still solve these relatively ill-conditioned problems, likely with no performance gain.
Table 5: Accuracy and iteration counts of MPGLS and MPGLS-GMRES-BD. ’er-1’, ’er-2’, and ’iter’ denote, respectively, $\lVert Wx+Vy-d\rVert_{2}/\bigl{(}\lVert W\rVert_{\mathsf{F}}\lVert x\rVert_{2}+\lVert V\rVert_{\mathsf{F}}\lVert y\rVert_{2}+\lVert d\rVert_{2}\bigr{)}$, $\big{\lvert}\lVert y\rVert_{2}/\lVert y_{\mathtt{dggglm}}\rVert_{2}-1\big{\rvert}$, which is related to $\kappa_{2}\bigl{(}[W,V]\bigr{)}$, and the number of iterations.
$\kappa_{2}\big{(}[W,V]\big{)}$ $\kappa=10^{3}$ $\kappa=10^{5}$ $\kappa=10^{7}$
MPGLS er-1 $2.0\cdot 10^{-17}$ $5.0\cdot 10^{-16}$ $9.4\cdot 10^{-15}$
er-2 $4.1\cdot 10^{-15}$ $1.0\cdot 10^{-11}$ $7.2\cdot 10^{-8}$
iter $3$ $4$ $13$
MPGLS-GMRES-BD er-1 $8.9\cdot 10^{-15}$ $2.9\cdot 10^{-14}$ $4.1\cdot 10^{-13}$
er-2 $2.3\cdot 10^{-15}$ $2.4\cdot 10^{-13}$ $8.7\cdot 10^{-12}$
iter $11$ $18$ $47$
@@ 7 Conclusions
Numerical experiments indicate that when the condition number $\kappa_{\infty}(\tilde{F})$ is not too large, our mixed precision algorithms enhance the performance significantly. Classical iterative refinement proves to be the fastest of the tested algorithms and only costs around 60% of the time of the LAPACK subroutines DGGLSE and DGGGLM. This shows that the mixed precision classical iterative refinement algorithm is preferably suitable in numerical algebra libraries for well-conditioned matrices. As the condition number increases, the classical iterative refinement algorithm may potentially fail to converge, and the mixed precision GMRES-based iterative refinement algorithm with the block-diagonal split preconditioner emerges as a favourable option. None of these iterative refinement based mixed precision algorithms achieves any acceleration for problems that are overly ill-conditioned.
@@
Theorem A.1.
and the following bounds are satisfied.
where $\eta_{0}$, $\eta_{1}$, and $\eta_{2}$ are small constants of $m$, $n$, $p$, and the unit roundoffs:
$\eta_{2}$ $=\max\bigl{\{}\epsilon_{\mathtt{trsv}}^{s}(n-p),\epsilon_{\mathtt{gemv}}^{s}(n-p),\epsilon_{\mathtt{gemv}}^{s}(m-n+p)\bigr{\}}.$
@@
$\operatorname{f{}l}(w_{1}-q_{1}-T_{12}y_{2})$ $=(I+\epsilon_{2}J_{2})\big{(}\hat{w}_{1}-\hat{q}_{1}-\operatorname{f{}l}(T_{12}y_{2})\big{)}$ (A.4)
@@
for $i=1$, $2$. The bounds of the vectors are subsequently summarized as
@@
Appendix B Rounding error regarding a “small” perturbation
@@
For the triangular system $Lx=v$ with $v$ containing a small perturbation, i.e., $Lx=(I+\epsilon J)v$, with $\lvert\epsilon\rvert\leq\beta<1$, $\lVert J\rVert_{2}=1$, there exists a matrix $\Delta L$ such that
@@
$\lVert\Delta L\rVert_{\mathsf{F}}\leq\frac{\beta+\epsilon_{\mathtt{trsv}}}{1-\beta}\lVert L\rVert_{\mathsf{F}}.$
@@
with $\lVert\Delta L_{1}\rVert_{\mathsf{F}}\leq\epsilon_{\mathtt{trsv}}\lVert L\rVert_{\mathsf{F}}$. Let $\Delta L_{2}=-\epsilon J(I+\epsilon J)^{-1}(L+\Delta L_{1})$. Then
@@
$\lVert\Delta L\rVert_{\mathsf{F}}$ $\leq\lVert\epsilon J(I+\epsilon J)^{-1}L\rVert_{\mathsf{F}}+\lVert(I+\epsilon J)^{-1}\Delta L_{1}\rVert_{\mathsf{F}}\leq\frac{\beta+\epsilon_{\mathtt{trsv}}}{1-\beta}\lVert L\rVert_{\mathsf{F}}.\qed$
@@
Let $H$ be a unitary matrix computed through Householder transformations. For applying $H$ to $v$ containing a small perturbation, i.e., $u=H\bigl{(}(I+\epsilon J)v\bigr{)}$, with $\lvert\epsilon\rvert\leq\beta<1$, $\lVert J\rVert_{\mathsf{F}}=1$,there exists $\delta u$ such that
@@
$\lVert\delta u\rVert_{2}\leq\frac{\beta+\epsilon_{\mathtt{unmqr}}}{1-\beta}\lVert\hat{u}\rVert_{2}.$
@@
where $\lVert\delta u_{1}\rVert_{2}\leq\epsilon_{\mathtt{unmqr}}\lVert\hat{u}\rVert_{2}$. Let $\delta u_{2}=\epsilon HJv$ and $\delta u=\delta u_{1}+\delta u_{2}$. Then $\hat{u}=Hv+\delta u$ with
$\lVert\delta u\rVert_{2}\leq\lVert\delta u_{1}\rVert_{2}+\lVert\delta u_{2}\rVert_{2}\leq\epsilon_{\mathtt{unmqr}}\lVert\hat{u}\rVert_{2}+\lvert\epsilon\rvert\cdot\lVert HJv\rVert_{2}.$
@@
$\lVert\delta u\rVert_{2}\leq\epsilon_{\mathtt{unmqr}}\lVert\hat{u}\rVert_{2}+\beta\bigl{(}\lVert\hat{u}\rVert_{2}+\lVert\delta u\rVert_{2}\bigr{)},$
@@
$\lVert\delta u\rVert_{2}\leq\frac{\beta+\epsilon_{\mathtt{unmqr}}}{1-\beta}\lVert\hat{u}\rVert_{2}.\qed$

### v3

@@ 5.3 Block-diagonal split preconditioner for the GLS problem
$\begin{split}M_{l}&=\operatorname{diag}\big{\{}\alpha^{-1/2}I_{p},\alpha^{1/2}\hat{T}_{2}^{-1}Q^{*},\alpha^{-1/2}\beta^{-1}S^{*}\hat{R}^{-*}\big{\}},\\ M_{r}&=\operatorname{diag}\big{\{}\alpha^{-1/2}I_{p},\alpha^{1/2}Q\hat{T}_{2}^{-*},\alpha^{-1/2}\beta^{-1}\hat{R}^{-1}S\big{\}},\end{split}$ (5.5)
@@
where $\check{I}_{m}=[I_{m},0]^{*}\in\mathbb{C}^{n\times m}$; $Z=[Z_{1}^{*},Z_{2}^{*}]^{*}$ with $Z_{2}\in\mathbb{C}^{p\times n}$. Note that the approximation pertains solely to the rounding error similarly to the arguments in Section 5.2. In exact arithmetic (5.6) is satisfied as an equality.
@@
$\begin{split}M_{l}&=\operatorname{diag}\bigl{\{}\alpha^{-1/2}I_{p},\alpha^{1/2}U^{-1}Q^{*},\alpha^{-1/2}\beta^{-1}Y^{*}\hat{R}^{-*}\bigr{\}},\\ M_{r}&=\operatorname{diag}\bigl{\{}\alpha^{-1/2}I_{p},\alpha^{1/2}QU^{-*},\alpha^{-1/2}\beta^{-1}\hat{R}^{-1}Y\bigr{\}},\end{split}$ (5.7)
@@
$\kappa_{\infty}(M_{l}\tilde{F}M_{r})\leq C\cdot(m+n+p),$
if $\mathcal{O}(\bm{u}_{f})\kappa_{2}(W)\kappa_{2}(V)\leq 1$.
[…]
@@
[…]
@@ 6 Numerical experiments
In this section, we present numerical results of our mixed precision LSE and GLS algorithms. In our tests, we set the combinations of precisions as ($\bm{u}_{f}$, $\bm{u}_{s}$, $\bm{u}$, $\bm{u}_{r}$) = (single, single, single, double) in IEEE standard. All experiments are performed on a Linux server equipped with two sixteen-core Intel Xeon Gold 6226R 2.90 GHz CPUs with 1024 GB of main memory.
@@ 6.1 Experiment settings
    DGGLSE [the reference algorithm]: the LAPACK subroutine using the nullspace approach via the GRQ factorization performed totally in double precision.
@@
    MPLSE-GMRES-Left: mixed precision LSE algorithm using GMRES with the left preconditioner (5.1) as iterative refinement.
@@
In our implementation, the GMRES algorithm is conducted in double precision entirely. Carson and Daužickaitė discussed using lower precision in GMRES-based iterative refinement in [9]. However, for dense matrices, there is no need for that, as the most time-consuming part, matrix–vector multiplication, needs to be performed in the working precision regardless. The parameter tol is set to $10^{-13}$ for the stopping criteria of iterative refinement.
@@ 6.2 Tests for the LSE problem
For the LSE problem, cases $m>n\gg p$ and $m\gg n>p$ are often of interest in practice. In case $m>n\gg p$, we choose $n\in\{1024,2048,3072\}$ in our experiments, and set $(m,p)=(8n,n/32)$, $(8n,n/64)$, $(10n,n/32)$, $(10n,n/64)$, $(12n,n/32)$, and $(12n,n/64)$ to generate different test matrices with fixed condition number $\kappa_{2}\big{(}[A^{*},B^{*}]\big{)}=10^{3}$, $10^{5}$, $10^{7}$, $10^{9}$, respectively. In case $m\gg n>p$, we choose $n\in\{128,192,256\}$, and set $(m,p)=(128n,n/8)$, $(128n,n/4)$, $(128n,n/2)$, $(256n,n/8)$, $(256n,n/4)$, and $(256n,n/2)$ with fixed condition number $\kappa_{2}\big{(}[A^{*},B^{*}]\big{)}=10^{5}$. The matrix $A^{*},B^{*}]$ with a specified condition number is generated through $A^{*},B^{*}]=W_{1}\Sigma W_{2}$, where $\Sigma$ is a diagonal matrix with the given condition number produced by DLATM1 from LAPACK, and $W_{1}$ and $W_{2}$ are the unitary factors from the QR factorization of randomly generated matrices. The parameters are set as $\alpha=\lVert r_{0}\rVert_{2}$ and $\beta=1$.
@@ 6.2.1 Case $m>n\gg p$
MPLSE err-1 $3.3\cdot 10^{-17}$ $0\cdot 10^{-16} 2.2$$2\cdot 10^{-14}$ $2.6\cdot 10^{-10}$
@@
Figure 1: Relative run time of MPLSE, MPLSE-GMRES-Left, and MPLSE-GMRES-BD compared to DGGLSE for matrices with $\kappa_{2}\bigl{(}[A^{*},B^{*}]\bigr{)}=10^{3}$. For each matrix, the three columns from left to right represent the results of MPLSE, MPLSE-GMRES-Left, and MPLSE-GMRES-BD, respectively.
Figure 2: Relative run time of MPLSE, MPLSE-GMRES-Left, and MPLSE-GMRES-BD compared to DGGLSE for matrices with $\kappa_{2}\bigl{(}[A^{*},B^{*}]\bigr{)}=10^{5}$. For each matrix, the three columns from left to right represent the results of MPLSE, MPLSE-GMRES-Left, and MPLSE-GMRES-BD, respectively.
Figure 3: Relative run time of MPLSE, MPLSE-GMRES-Left, and MPLSE-GMRES-BD compared to DGGLSE for matrices with $\kappa_{2}\bigl{(}[A^{*},B^{*}]\bigr{)}=10^{7}$. For each matrix, the three columns from left to right represent the results of MPLSE, MPLSE-GMRES-Left, and MPLSE-GMRES-BD, respectively.
From Table 4 and Figures 13, we observe that in comparison with the left preconditioner, the GMRES-based algorithm with the block-diagonal split preconditioner requires more iterations to converge. However, the left preconditioner costs more execution time due to the higher computational cost of applying (5.1) to a vector. This indicates that it is preferable to use the block-diagonal split preconditioner (5.3) or (5.4) for GMRES-based iterative refinement.
@@
As the condition number $\kappa_{2}\bigl{(}[A^{*},B^{*}]\bigr{)}$ becomes $10^{7}$, which is close to $\bm{u}_{f}^{-1}$, all three mixed precision LSE algorithms require more iterations and longer time to converge; see Figure 3. This is due to the reduced accuracy of the initial guess and GRQ factorization, which also affects the quality of the preconditioner. Our analysis indicates that MPLSE is not suitable for scenarios where condition numbers exceed approximately $\bm{u}_{f}^{-1}$. As demonstrated by the experiments in Table 4 with $\kappa=10^{9}$, MPLSE fails to converge in this case. In Figure 3, we see that MPLSE-GMRES-BD remains approximately $10\%$$20\%$ faster than DGGLSE, a speedup between 1.11 and 1.25. However, that will diminish rapidly as the condition number increases further.
@@ 6.2.2 Case $m\gg n>p$
Figure 4: Relative run time of MPLSE, MPLSE-GMRES-Left, and MPLSE-GMRES-BD compared to DGGLSE for matrices with $\kappa_{2}\bigl{(}[A^{*},B^{*}]\bigr{)}=10^{5}$. For each matrix, the three columns from left to right represent the results of MPLSE, MPLSE-GMRES-Left, and MPLSE-GMRES-BD, respectively.
@@ 6.3 Tests for the GLS problem
For the GLS problem, we choose the dimension $n\in\{1024,2048,3072\}$ in our experiments, and set $(p,m)=(8n,n/32)$, $(8n,n/64)$, $(10n,n/32)$, $(10n,n/64)$, $(12n,n/32)$, and $(12n,n/64)$ to generate different test matrices with fixed condition number $\kappa_{2}\bigl{(}[W,V]\bigr{)}=10^{3}$, $10^{5}$, $10^{7}$, $10^{9}$, respectively. Note that we do not test GMRES with the left preconditioner (5.2) for the GLS problem, since it has already been shown in Section 6.2 that it is preferable to use the block-diagonal split preconditioner. Similarly, the parameters are set as $\alpha=\lVert y_{0}\rVert_{2}$ and $\beta=1$.
@@
Table 5 presents the accuracy of the results and the iteration counts for a single test matrix of consistent size across varying condition numbers. In Figures 57, we present the relative run time for matrices with fixed condition numbers $\kappa_{2}\bigl{(}[W,V]\bigr{)}=10^{3}$, $10^{5}$, $10^{7}$. For test matrices with relatively small condition numbers, both algorithms reduce the computational time by a factor around $40\%$$50\%$, i.e., a speedup between $1.66$ and $2.00$, with MPGLS being slightly faster. The analysis shows that MPGLS fails when condition numbers exceed approximately $\bm{u}_{f}^{-1}$, which is also confirmed by the results presented in Table 5 with $\kappa=10^{9}$. As demonstrated in Table 5, MPGLS-GMRES-BD can still solve these relatively ill-conditioned problems, likely with no performance gain.
Table 5: Accuracy and iteration counts of MPGLS and MPGLS-GMRES-BD. ‘err-1’, ‘err-2’, and ‘iter’ denote, respectively, $\lVert Wx+Vy-d\rVert_{2}/\bigl{(}\lVert W\rVert_{\mathsf{F}}\lVert x\rVert_{2}+\lVert V\rVert_{\mathsf{F}}\lVert y\rVert_{2}+\lVert d\rVert_{2}\bigr{)}$, $\big{\lvert}\lVert y\rVert_{2}/\lVert y_{\mathtt{dggglm}}\rVert_{2}-1\big{\rvert}$, which is related to $\kappa_{2}\bigl{(}[W,V]\bigr{)}$, and the number of iterations.
$\kappa_{2}\big{(}[W,V]\big{)}$ $\kappa=10^{3}$ $\kappa=10^{5}$ $\kappa=10^{7}$ $\kappa=10^{9}$
MPGLS err-1 $2.0\cdot 10^{-17}$ $5.0\cdot 10^{-16}$ $9.4\cdot 10^{-15}$ $8.7\cdot 10^{-10}$
err-2 $4.1\cdot 10^{-15}$ $1.0\cdot 10^{-11}$ $7.2\cdot 10^{-8}$ $6.8\cdot 10^{-1}$
iter 33 44 $13$ diverged
MPGLS-GMRES-BD err-1 $8.9\cdot 10^{-15}$ $2.9\cdot 10^{-14}$ $4.1\cdot 10^{-13}$ $6.7\cdot 10^{-11}$
err-2 $2.3\cdot 10^{-15}$ $2.4\cdot 10^{-13}$ $8.7\cdot 10^{-12}$ $1.1\cdot 10^{-8}$
iter $11$ $18$ $47$ $297$
@@ 7 Conclusions
[…]
@@
$\operatorname{f{}l}(w_{1}-q_{1}-T_{12}y_{2})$ $=(I+\epsilon_{2}J_{2})\big{(}\hat{w}_{1}-\hat{q}_{1}-\operatorname{f{}l}(T_{12}y_{2})\big{)},$ (A.4)
@@
for $i=1$, 22. The bounds of the vectors are subsequently summarized as
@@
Appendix B Rounding error regarding a ‘small’ perturbation
@@
Let $L\in\mathbb{C}^{n\times n}$, $v\in\mathbb{C}^{n}$. For the triangular system $Lx=v$ with $v$ containing a small perturbation, i.e., $Lx=(I+\epsilon J)v$, with $\lvert\epsilon\rvert\leq\beta<1$, $\lVert J\rVert_{2}=1$, there exists a matrix $\Delta L$ such that
@@
$\lVert\Delta L\rVert_{\mathsf{F}}\leq\frac{\beta+\epsilon_{\mathtt{trsv}}(n)}{1-\beta}\lVert L\rVert_{\mathsf{F}}.$
@@
with $\lVert\Delta L_{1}\rVert_{\mathsf{F}}\leq\epsilon_{\mathtt{trsv}}(n)\cdot\lVert L\rVert_{\mathsf{F}}$. Let $\Delta L_{2}=-\epsilon J(I+\epsilon J)^{-1}(L+\Delta L_{1})$. Then
@@
$\lVert\Delta L\rVert_{\mathsf{F}}$ $\leq\lVert\epsilon J(I+\epsilon J)^{-1}L\rVert_{\mathsf{F}}+\lVert(I+\epsilon J)^{-1}\Delta L_{1}\rVert_{\mathsf{F}}\leq\frac{\beta+\epsilon_{\mathtt{trsv}}(n)}{1-\beta}\lVert L\rVert_{\mathsf{F}}.\qed$
@@
Let $H\in\mathbb{C}^{n\times n}$ be a unitary matrix computed through Householder transformations, $v\in\mathbb{C}^{n}$. For applying $H$ to $v$ containing a small perturbation, i.e., $u=H\bigl{(}(I+\epsilon J)v\bigr{)}$, with $\lvert\epsilon\rvert\leq\beta<1$, $\lVert J\rVert_{\mathsf{F}}=1$, there exists $\delta u$ such that
@@
$\lVert\delta u\rVert_{2}\leq\frac{\beta+\epsilon_{\mathtt{unmqr}}(n)}{1-\beta}\lVert\hat{u}\rVert_{2}.$
@@
where $\lVert\delta u_{1}\rVert_{2}\leq\epsilon_{\mathtt{unmqr}}(n)\cdot\lVert\hat{u}\rVert_{2}$. Let $\delta u_{2}=\epsilon HJv$ and $\delta u=\delta u_{1}+\delta u_{2}$. Then $\hat{u}=Hv+\delta u$ with
$\lVert\delta u\rVert_{2}\leq\lVert\delta u_{1}\rVert_{2}+\lVert\delta u_{2}\rVert_{2}\leq\epsilon_{\mathtt{unmqr}}(n)\cdot\lVert\hat{u}\rVert_{2}+\lvert\epsilon\rvert\cdot\lVert HJv\rVert_{2}.$
@@
$\lVert\delta u\rVert_{2}\leq\epsilon_{\mathtt{unmqr}}(n)\cdot\lVert\hat{u}\rVert_{2}+\beta\bigl{(}\lVert\hat{u}\rVert_{2}+\lVert\delta u\rVert_{2}\bigr{)},$
@@
$\lVert\delta u\rVert_{2}\leq\frac{\beta+\epsilon_{\mathtt{unmqr}}(n)}{1-\beta}\lVert\hat{u}\rVert_{2}.\qed$